OpenStudy (anonymous):

What is 15 divided by the square root of 7, in simplied form?

OpenStudy (anonymous):

15 √7 / 7

OpenStudy (phi):

to do this problem, multiply top and bottom by sqrt(7) like this \[ \frac{15}{\sqrt{7}}\cdot \frac{\sqrt{7}}{\sqrt{7}}\]